OpenStudy (anonymous):

verify that (sinxcosx+cosx)/(sinx+sin^2x)=cotx

zepdrix (zepdrix):

\[\Large \frac{\sin x \cos x + \cos x}{\sin x + \sin^2x}\quad=\]Factoring a cosine out of each top term, and a sine out of each bottom term gives us, \[\Large \frac{\cos x(\sin x+1)}{\sin x(1+\sin x)}\quad=\]

zepdrix (zepdrix):

Do you see the cancellation we can make from this spot?

OpenStudy (anonymous):

(sinxcosx+cosx)/(sinx+sin^2x) =cosx(sinx+1)/sinx(1+sinx) =cosx/sinx =cotx

OpenStudy (anonymous):

yes! we can cancel out the (sinx-1) so that we get \[\frac{ \cos x }{ \sin x } \]

zepdrix (zepdrix):

yes good :)